It takes a boy 90 minutes to mow the lawn, but his sister can mow it in 60 minutes. How long would it take them to mow the lawn if they worked together, using two lawn mowers?
Answer by ptaylor(2198)

Let x=time required to mow the lawn if both work together
Then both work at the rate of 1/x lawn per min
boy works at the rate of 1/90 lawn per min
sis works at the rate of 1/60 lawn per min
Then both work at the rate of 1/90 + 1/60=2/180+3/180=5/180=1/36 lawn per min
So our equation to solve is
1/36=1/x multiply each side by 36x
x=36 min----time it takes both working together
CK
in 36 min boy mows (1/90)*36=36/90=6/15 of the lawn
in 36 min girl mows (1/60)*36 =36/60 =9/15 of the lawn
6/15 + 9/15=15/15=1 (lawn that is)
Hope this helps---ptaylor